Circular No. 8322

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 EW_{38} (CATALINA-LINEAR) An apparently asteroidal object reported independently by the Cata-

lina and LINEAR surveys (discovery observations below from MPS 102307; linked by G. V. Williams) has been found to be cometary on CCD images obtained at two Arizona observatories. Observations (via independent discovery) on Apr. 14.3 UT by M. T. Read and J. A. Larsen, using the 0.9-m f/3 Spacewatch reflector at Kitt Peak, show a tail 100′′ long in p.a. 300^{o}, extending asymmetrically from the south part of the nuclear condensation. Exposures taken to look for cometary appearance on Apr. 14.32 by C. W. Hergenrother with the 1.2-m reflector at Mt. Hopkins show a very condensed 9′′ coma and a narrow tail 210′′ long in p.a. 295^{o}.

2004 UT  α_{2000}  δ_{2000}  Mag. Observer
Mar. 14.46190 13^{h}57^{m}45.^{s}77 −4^{o}38′44′′.6 20.0 Catalina
16.33976 13 56 53.63 −4 37 59.8 19.7 LINEAR

Additional astrometry, the following elliptical orbital elements, and an ephemeris appear on MPEC 2004-G46.

Epoch = 2003 Nov. 17.0 TT

T = 2003 Nov. 19.8968 TT  ω = 90.^{o}2353
e = 0.500268  Ω = 49.9014 }2000.0
q = 1.793350 AU  i = 6.5233

a = 3.588624 AU n^{o} = 0.1449814 P = 6.798 years

SUPERNOVAE 2004bc AND 2004be Further to IAUC 8321, T. Matheson et al. report that a spectrum of

SN 2004bc (cf. IAUC 8315), obtained by J. Huchra on Apr. 12.26 UT with the Mt. Hopkins 1.5-m telescope, shows it to be a type-Ia supernova, a few days before maximum. Adopting the NED recession velocity of 7221 km/s for the host galaxy, the supernova expansion velocity is 11700 km/s for Si II (rest 635.5 nm). The spectrum is highly reddened and shows a strong narrow Na I D absorption at the velocity of the host galaxy with an equivalent width of 0.2 nm. The dust maps of Schlegel et al. (1998, Ap.J. 500, 525) indicate a low value for Galactic reddening of E(B − V) = 0.065. A spectrogram of SN 2004be (cf. IAUC 8317), obtained by Huchra on Apr. 13.23, shows it to be a type-II supernova; adopting the NED recession velocity of 2282 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 9800 km/s.

2004 April 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green